THE THZ SPECTRUM OF GLYCOLALDEHYDE

\underline{MANUEL GOUBET}, THERESE R. HUET, IMANE HAYKAL, LAURENT MARGULES, Laboratoire PhLAM, UMR8523 CNRS-Universite Lille 1, F-59655 Villeneuve d’Ascq Cedex, France; OLIVIER PIRALI, PASCALE ROY, Ligne AILES - Synchrotron SOLEIL, L’Orme des Merisiers Saint Aubin, F-91192 Gif-sur-Yvette, France.

The vibration-rotation spectrum of the ν_{1}-0, ν_{2}-0 and ν_{3}-0 bands of glycolaldehyde was recorded up to 12 THz, using the far-infrared beamline AILES at the synchrotron SOLEIL and a Fourier transform spectrometer coupled to a multipass cell. More than eight thousands lines were assigned, revealing the rotation structure up to J=80, K_{a}=38 for the ground state. The THz data were fitted simultaneously with pure rotational transitions of better accuracy observed in the microwave (1), in the millimeter-wave (2) and in the sub-millimeter-wave (3) range. In addition new data were recorded at Lille in the 150-300 GHz and 750-950 GHz range. The THz lines and the microwave - (sub)-millimeterwave lines are reproduced with a standard deviation of 2 10^{−4} cm^{−1} and 40 KHz, respectively. Glycolaldehyde has been identified toward the galactic center (4). The vibrational state partition function can be reevaluated according to the bands origins associated with ν_{1}, ν_{2}, and ν_{3}, which are observed experimentally for the first time. This work is supported by the Programme National de Physico-Chimie du Milieu Interstellaire (PCMI-CNRS) and by the contract ANR-08-BLAN-0054.

1. M. Rey, J.-R. Aviles-Moreno and T. R. Huet, Chem. Phys. Lett. 430(2006) 121 ; K.-M. Marstokk and H. Mollendal, J. Mol. Struct. 5 (1970) 205.

2. R. A. H. Butler, F. C. De Lucia, D. T. Petkie, H. Mollendal, A. Horn, and E. Herbst, ApJS 134 (2001) 319. ; S. L. Widicus-Weaver, R. A. H. Butler, B. J. Drouin, D. T. Petkie, K. A. Dyl, F. C. De Lucia, and G. A. Blake, ApJ 158(2005)188.

3. P. B. Carroll, B. J. Drouin, and S. L. Widicus-Weaver, ApJ 723 (2010) 845.

4. J. M. Hollis, S. N. Vogel, L. E. Snyder, P. R. Jewell, and F. J. Lovas, ApJ 554 (2001) L81. ; M.T. Beltran, C. Codella, S. Viti, R. Niri, R. Cesaroni, ApJ 690 (2009) L93.